OpenStudy (anonymous):

Physics Question: Mechanics Circular Motion A stone is rotated steadily in a horizontal circle with a period T by a string of length l. if the tension in the string is constant and l increases by 1% find the percentage change in T.

OpenStudy (anonymous):

The answer should 0.5%

OpenStudy (perl):

in circular motion, F = mv^2/r

OpenStudy (perl):

woops , T = 2pi*r/v

OpenStudy (perl):

here the centripetal force is provided by the tension in the string, so Tension = m*v^2/r

OpenStudy (anonymous):

ok do i have to use ratios at any time?

OpenStudy (perl):

you dont have to, if you can find another way

OpenStudy (anonymous):

can you show me how to solve it then please thanks

OpenStudy (perl):

since Force is constant we have

OpenStudy (perl):

m*v^2/ r = m *x^2 / (1.01*r) , because the radius increased by 1% , solve for x, the new speed

OpenStudy (anonymous):

ok

OpenStudy (perl):

did you get x = sqrt(1.01)*v ?

OpenStudy (anonymous):

yep

OpenStudy (perl):

ok great, now plug this into the T (period) equation

OpenStudy (perl):

T = 2pi * r / v

OpenStudy (perl):

plug this 'new' speed in for v

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (perl):

Let's call the new period T '. The old period, with the old radius and old speed was T = 2pi * r / v The new speed is sqrt(1.01)*v The new radius is 1.01*r Plugging in we have T ' = 2*pi * (1.01*r) / ( sqrt(1.01)* v ) = (2pi * r /v ) * 1.01 / sqrt( 1.01) = T * 1.01 / sqrt(1.01) = T * 1.004987 = T ( 1 + .005) approximately = T( 1 + .5% )

OpenStudy (perl):

@Michele_Laino can you double check this is correct

OpenStudy (anonymous):

i was copying your earlier posts and some seem to have been deleted

OpenStudy (perl):

the crucial point to notice here is that if F = mv^2 / r , and F is constant (since F is tension and it is given that tension is constant), then if you increase radius then you must increase velocity in order for the fraction to stay the same. Think about a fraction, if i want to increase the bottom, i have to increase the top as well if i want keep the fraction the same number overall

OpenStudy (perl):

or you could increase mass, but in the problem mass is fixed , so that forces velocity to change

OpenStudy (anonymous):

yes thats clear

OpenStudy (anonymous):

ok

OpenStudy (perl):

so what we had initially was the equation: Tension = mv^2 / r = m v ' ^2 / r ' , since tension stayed constant. we were given that r ' = 1.01 * r and we solved for v ' in terms of v

OpenStudy (perl):

and you see how we got the period T equation ( T does not stand for tension but for time of period)

OpenStudy (perl):

we know that average speed = distance / time , In a circle, since we have constant speed we can choose distance to be a one revolution , and T is the time it takes to make one revolution (also known as period). v = circumference / T v = 2pi*r / T solve for T T = 2pi * r / v

OpenStudy (perl):

I hope that helped :)

OpenStudy (anonymous):

it did clearest explanation i have had so far thanks a lot

OpenStudy (perl):

also you should check to see if the answer makes intuitive sense. The answer we got is that the period increased by 0.5% when we increased radius by 1% but kept tension fixed. Does this make sense? Yes because there will be a larger circumference to cover so it makes sense that period (time to complete 1 revolution) increases.